OpenStudy (anonymous):

Complete the standard form of the equation y=10-5/2 x

OpenStudy (amistre64):

well standard form has a definition to it. can you tell me what standard form is?

OpenStudy (amistre64):

in your material, standard form should look like: ax + by = c notice that we have both x and y parts on one side .... what should we do to work this?

OpenStudy (amistre64):

kind of, we can subtract -5/2 x from each side, but this can also be called addition ... add 5/2 x to each side

OpenStudy (amistre64):

y = 10 - 5/2 x +5/2x +5/2 x ------------------ 5/2 x + y = 10 almost there ....

OpenStudy (amistre64):

we want all integer values in this, not fraction or decimals ... so we need to make 5/2 into just a 5 any thoughts on how to do this?

OpenStudy (amistre64):

well, multiplyng everything by 2 should work for us 2 (5/2 x + y = 10) 5*2/2 x + 2y = 2(10) since 2/2 = 1 ... 5x + 2y = 20 now we need to make sure the leading x term is positive ... in this case it is so we are fine

OpenStudy (amistre64):

yeah, you editing out the content of your posts does not constitute me breaking the rules ... good luck with that tho

OpenStudy (e.mccormick):

See, mods can see deleted posts, so what you did by "deleting" basically means nothing to a mod.